OpenStudy (anonymous):

Differentiation and integration for y=2cos 3x/e^{2x}

OpenStudy (anonymous):

y'=(2.3(-sin3x).e^2x-2cos3x.2.e^2x)/(e^2x)^2 =(-6sin(3x)e^2x-4cos(3x)e^2x)/e^4x

OpenStudy (anonymous):

-6e^(2x)*sin3x-4cos3x*e^(2x) all over e^(4x)

OpenStudy (anonymous):

iono integrate it...

OpenStudy (anonymous):

integration solution is long if you want,i will study write it.

OpenStudy (anonymous):

No,thank you cnsu 90.

OpenStudy (anonymous):

cos3x=u -3sin3xdx=du and e^-2x dx=dv v= -2e^2x I=uv-int_vdu then integrate 2 times